OpenStudy (anonymous):

solve : log base 4 of (x^2 - 3) + log base 4 of 10=1

OpenStudy (zzr0ck3r):

log(a) + log(b) = log(a*b) so log_4(10x^2 - 30) = 1 so 4^1 = 10x^2 - 30 34/10 = x^2 x = +-sqrt(17/5)